# Nash Equilibria in Games with Playerwise Concave Coupling Constraints: Existence and Computation

### v2

@@ C.3 Berge’s maximum theorem
Next, we provide a statement of Berge’s maximum theorem that is used in our existence proof. This particular variant considers maximization of strongly concave parameterized functions over a parameterized constraint set. For reference, see for instance the version stated in Sundaram, (1996), Theorem 9.17.
Theorem 1.
@@ C.4 Fixed point theorem
As pointed out by Debreu, (1952), in finite dimension, Begle’s fixed point theorem applies to compact metric spaces that are both contractible and locally contractible. Local contractibility is a property of the space’s local geometry that holds for any subset of Euclidean space (see notes by Wilkins, (2017)). Hence, it suffices to assume contractibility. The statement then follows from Begle, (1950), Theorem 1. Note that we state the fact for single-valued functions which directly follows from the set-valued version. ∎
@@ Appendix D Proof of Existence Result
Lemma 1.
Suppose Assumption 1 holds. Let $\mathcal{C}\not$$\emptyset$, and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be a connected component of $\mathcal{C}$. For any $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, the set
@@
Lemma 2.
Suppose Assumption 2 holds. Let $\mathcal{C}\not$$\emptyset$, and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be a connected component of $\mathcal{C}$. Then, for all $i\in[m]$, the set-valued function $\mathcal{C}^{\prime}_{i}(\cdot)$ is continuous (i.e. both upper and lower semicontinuous) over $\operatorname{supp}(\mathcal{C}^{\prime}_{i})$.
@@
We apply Lemma 6.1 of Still, (2018)’s notes on parametric optimization. In our setting, it states that if the MFCQ condition holds at every $x\in\mathcal{C}^{\prime}$ (which is given by our Assumption 2), then the parametric set-valued function $\mathcal{C}^{\prime}_{i}(\cdot)$ is $L$-Lipschitz continuous in terms of the Hausdorff distance over $\operatorname{supp}(\mathcal{C}^{\prime}_{i})$ for large enough $L>0$. It is well-known that for closed-valued functions such Hausdorff continuity implies upper and lower semicontinuity (see e.g. Clarke et al., (1998), 7.29). ∎
@@
Regarding (a), observe that each $h_{i}$ is well-defined: By definition, we have $x_{i}\in\mathcal{C}^{\prime}_{i}(x_{-i})$. By convexity of $\mathcal{C}^{\prime}_{i}(x_{-i})$, as shown in Lemma 1, it holds that $\overline{c}(\mathcal{C}^{\prime}_{i}(x_{-i}))\in\mathcal{C}^{\prime}_{i}(x_{-i})$. Thus for all $t\in[0,1]$, again by convexity, $(1-t)x_{i}+t\,\overline{c}(\mathcal{C}^{\prime}_{i}(x_{-i}))\in\mathcal{C}^{\prime}$. As $h_{i}$ keeps $x_{-i}$ fixed, we have $h_{i}(t,x)\in\mathcal{C}^{\prime}$ for all $t\in[0,1]$ and $x\in\mathcal{C}^{\prime}$. Then, by definition, $H_{1}$ is well-defined, and the same follows for each $H_{i}$ inductively. We conclude that $H$ is well-defined, as it is piecewise equivalent to the $H_{i}$’s.
To show (b), note that by Lemma 2, $\mathcal{C}^{\prime}_{i}(x_{-i})$ is continuous in $x_{-i}$. Then, continuity of each $h_{i}$ in $(t,x)$ follows together with the continuity of the centroid function $\overline{c}(\cdot)$. Similarly as in (a), the recursive definitions of each $H_{i}$ and the piecewise definition of $H$ lets us conclude that $H$ is indeed continuous in $(t,x)$. In particular, we point out that continuity at the boundary of the pieces is ensured, as for all $i\in[m]\setminus\left\{1\right\}$ and all $x\in\mathcal{C}^{\prime}$, by definition, $\lim_{t\to{(i-1)/m}^{+}}H_{i}(t,x)=H_{i-1}(\varphi^{-1}_{i-1}(1),x)$.
@@
By a variant of Berge’s maximum theorem for strongly concave objective functions over convex constraints (see Theorem 1 in Appendix C), $\Psi_{i}$ is continuous and single-valued. Define the joint best response map $\Psi:\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ as the composition $\Psi\coloneqq\Psi_{1}\circ\dots\circ\Psi_{m}$, which is also continuous.
@@ E.2 Technical lemma
We provide a lemma which is applied in the proof of Lemma 11 in Appendix F. Informally, it says the following: If we perform a projected gradient ascent step along a (locally) smooth function $f$ and the observed gradients of $f$ are perturbed by some error $\delta$, then, as long as $\delta$ is sufficiently small, the value of $f$ will not decrease.
Lemma 4.
@@
We restate the following sufficient increase inequality for projected gradient ascent that will be used in the proof of Theorem 2. The claimed inequality is standard; for reference, see Bubeck et al., (2015), Lemma 3.6.
Lemma 5.
@@ F.1 Mixed-strategy potential games
In Example 3 we claimed that when considering mixed strategies over a finite action potential game, the potential structure is lifted to the continuous action game over compact convex set of mixed strategies. Here we state the corresponding proposition.
Proposition 2.
@@
For the case of unconstrained potential games, Proposition 2 is a direct consequence of Monderer and Shapley, (1996), Lemma 2.10. The same reasoning extends straightforwardly to the constrained case; we therefore omit the proof here.
@@ F.2 Technical lemmas
Lemma 7.
Let $\Phi:\mathcal{X}\to\mathbb{R}$ be a potential function for our game with utilities $\{u_{i}\}_{i\in[m]}$. Then $\Phi$ is $(M\sqrt{m})$-smooth in the joint strategy $x\in\mathcal{X}$, where $M$ is the smoothness parameter of the utilities, see Assumption 4.
@@
Note that differentiability of $u_{i}$ for all $i\in[m]$ (Assumption 4) implies differentiability of $\Phi$, and it is well known that for all $x\in\mathcal{X}$, $\nabla_{x_{i}}\Phi(x)=\nabla_{x_{i}}u_{i}(x)$. Using the smoothness of $u_{i}$ from Assumption 4, we get
@@ F.3 Smoothness along the trajectory
where we use $L$-Lipschitz continuity of $u_{i}$ and $c_{j}$, see Assumption 4, for bounding the gradient norms. Then, we have for each $j\in[b]$,
@@
Using $M$-smoothness and $L$-Lipschitz continuity of $c_{j}$ for all $j\in[b]$, as well as $(M\sqrt{m})$-smoothness of $\Phi$, as shown in Lemma 7, we then bound the Hessian norm as follows:
@@
Lemma 10.
Suppose $x^{(t)}\in\mathcal{C}$ is $\beta$-feasible and each $\gamma^{(t)}$ is bounded as in the statement of Lemma 8. Then $\Phi^{\eta}$ is $M^{\eta,\beta/2}_{\Phi}$-smooth over $x^{(t)},x^{(t+1)}]$ where $M^{\eta,\beta}_{\Phi}$ is defined as in Lemma 9, i.e. for all $x,x^{\prime}\in[x^{(t)},x^{(t+1)}]$,
@@
Taking the norm, and using the fact that $x^{\prime}+t(x-x^{\prime})$ is $\beta/2$-feasible for all $t\in[0,1]$ as shown in Lemma 8, together with the Hessian norm bound from Lemma 9, we get
@@ F.4 Lower bounding the stepsize
Motivated by the bound in Lemma 8 and smoothness along the trajectory as shown in Lemma 10, we choose the stepsize at iteration $t$ as
@@
In this subsection, we want to argue that the above defined stepsizes can be uniformly lower bounded by a constant multiple of $\eta$. This is necessary to guarantee sufficient progress in our convergence analysis. We start by showing that feasibility margins do not decrease further once a point sufficiently close to the boundary is reached. Based on Lemma 4, we prove the following statement. Note that a similar statement was proven in Usmanova et al., (2024), Fact 1, with the difference that ours is for projected gradient steps, which requires a different approach.
@@
Lemma 11.
Let Assumption 5 hold with $\rho\geq\eta$. If at any iteration $0\leq t\leq T-1$, we have $\beta^{(t)}\leq\overline{c}\eta$ with $\overline{c}=\frac{\ell}{6L(b+1)}$, then for iteration $t+1$, we get $\prod_{j\in\mathcal{J}}\beta_{j}^{(t+1)}\geq\prod_{j\in\mathcal{J}}\beta_{j}^{(t)}$ for any $\mathcal{J}$ such that $\mathcal{J}_{t}\subseteq\mathcal{J}$ with $\mathcal{J}_{t}\coloneqq\{j\in[b]\mid\beta_{j}^{(t)}\leq\eta\}$.
@@
We aim to apply Lemma 4 with $x^{(t)}, x+=x$$^{+}=x^{(t+1)}$,
@@
Note that by our choice of $\gamma^{(t)}$, by Lemma 10, we have $\lVert\nabla f(x^{+})-\nabla f(x)\rVert\leq\frac{1}{\gamma^{(t)}}\lVert x^{+}-x\rVert$. In order to apply Lemma 4, we further need to show that for $F(x)\coloneqq\frac{1}{\gamma^{(t)}}\left(x^{+}-x\right)$, it holds that $\lVert F(x)\rVert\geq 3\lVert\delta\rVert$.
We begin by giving a lower bound on $\lVert F(x^{(t)})\rVert$. By Proposition B.1 of Agarwal et al., (2021) and Assumption 5, we have
@@
where $(\star)$ is due to $\beta_{j}^{(t)}>\eta$ for $j\not\in\mathcal{J}_{t}$. Combining the two bounds, and by definition of $\overline{c}$, we obtain from Lemma 4 that
@@
As a consequence of Lemma 11, we obtain the following lower bound on the distance of iterates to the boundary of $\mathcal{C}$. Having established Lemma 11 above, the proof of Lemma 12 is equivalent to the proof of an analogous result stated as Lemma 6 in Usmanova et al., (2024); we omit the proof here.
Lemma 12.
Let Assumptions 3 and 5 hold with $\rho\geq\eta$. Then, for all $t\geq 1$, it holds that $\beta^{(t)}\geq c\eta$ where $c\coloneqq\frac{1}{2}\left(\frac{\ell}{24L(b+1)}\right)^{b}$.
@@
First, we argue that our independent updates are equivalent to centralized gradient ascent on the regularized potential function $\Phi^{\eta}$. This follows from the fact that, by Lemma 6, the regularized utilities form a potential game, meaning we have $\nabla_{x_{i}}\Phi^{\eta}(x)=\nabla_{x_{i}}B^{\eta}_{i}(x)$ for any $x\in\mathcal{X}$ and $i\in[m]$. Together with separability of the projection operator onto the Cartesian product of the sets $\mathcal{X}_{i}$, see Leonardos et al., (2022), Lemma D.1, we obtain
@@
By smoothness along the trajectory and our choice of stepsizes, by Lemma 5, for all $0\leq t\leq T-1$,
@@
Together with the fact that there exists a constant $C>0$ such that $\gamma^{(t)}\geq C\eta$, see Lemma 13, and by telescoping the sum, we obtain
@@
We note that using Lemma 12, $\Phi^{\eta}$ can be bounded by a constant. Thus for $\eta=\epsilon$ and $T=\mathcal{O}(\epsilon^{-3})$, there exists $0\leq t^{\star}\leq T-1$ such that $\lVert x^{(t^{\star}+1)}-x^{(t^{\star})}\rVert\leq\frac{\epsilon}{4}$. Next, we argue that then $x^{(t^{\star}+1)}$ satisfies the $\frac{\epsilon}{2}$-KKT conditions as introduced in Appendix E. Consider the pair of primal and dual variables given by
@@
    Primal feasibility: Lemma 8 implies that all iterates remain within $\mathcal{C}$. In particular, we have $x^{\star}\in\mathcal{C}$.
@@
    Lagrangian stationarity: By a standard property of projected gradient ascent, see e.g. Agarwal et al., (2021), Proposition B.1, $\lVert x^{(t^{\star}+1)}-x^{(t^{\star})}\rVert\leq\frac{\epsilon}{4}$ implies that
@@
Finally, Lemma 3 implies that the $\frac{\epsilon}{2}$-KKT strategy $x^{\star}$ constitutes an $\epsilon$-approximate constrained Nash equilibrium. ∎

### v3

@@ C.3 Berge’s Maximum Theorem
Next, we provide a statement of Berge’s maximum theorem that is used in our existence proof. This particular variant considers maximization of strongly concave parameterized functions over a parameterized constraint set. For reference, see for instance the version stated in Sundaram (1996), Theorem 9.17.
Theorem C.3.
@@ C.4 Fixed Point Theorem
As pointed out by Debreu (1952), in finite dimension, Begle’s fixed point theorem applies to compact metric spaces that are both contractible and locally contractible. Local contractibility is a property of the space’s local geometry that holds for any subset of Euclidean space (see notes by Wilkins (2017)). Hence, it suffices to assume contractibility. The statement then follows from Begle (1950), Theorem 1. Note that we state the fact for single-valued functions which directly follows from the set-valued version. ∎
@@ Appendix D Proof of Existence Result
Lemma D.1.
Suppose Assumption 2.2 holds. Let $\mathcal{C}\not$$\emptyset$, and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be a connected component of $\mathcal{C}$. For any $i\in[m]$ and $x_{-i}\in\mathcal{X}_{-i}$, the set
@@
Lemma D.2.
Suppose Assumption 3.1 holds. Let $\mathcal{C}\not$$\emptyset$, and let $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ be a connected component of $\mathcal{C}$. Then, for all $i\in[m]$, the set-valued function $\mathcal{C}^{\prime}_{i}(\cdot)$ is continuous (i.e. both upper and lower semicontinuous) over $\operatorname{supp}(\mathcal{C}^{\prime}_{i})$.
@@
We apply Lemma 6.1 of Still (2018)’s notes on parametric optimization. In our setting, it states that if the MFCQ condition holds at every $x\in\mathcal{C}^{\prime}$ (which is given by our Assumption 3.1), then the parametric set-valued function $\mathcal{C}^{\prime}_{i}(\cdot)$ is $L$-Lipschitz continuous in terms of the Hausdorff distance over $\operatorname{supp}(\mathcal{C}^{\prime}_{i})$ for large enough $L>0$. It is well-known that for closed-valued functions such Hausdorff continuity implies upper and lower semicontinuity (see e.g. Clarke et al. (1998), 7.29). ∎
@@
Regarding (a), observe that each $h_{i}$ is well-defined: By definition, we have $x_{i}\in\mathcal{C}^{\prime}_{i}(x_{-i})$. By convexity of $\mathcal{C}^{\prime}_{i}(x_{-i})$, as shown in Lemma D.1, it holds that $\overline{c}(\mathcal{C}^{\prime}_{i}(x_{-i}))\in\mathcal{C}^{\prime}_{i}(x_{-i})$. Thus for all $t\in[0,1]$, again by convexity, $(1-t)x_{i}+t\,\overline{c}(\mathcal{C}^{\prime}_{i}(x_{-i}))\in\mathcal{C}^{\prime}$. As $h_{i}$ keeps $x_{-i}$ fixed, we have $h_{i}(t,x)\in\mathcal{C}^{\prime}$ for all $t\in[0,1]$ and $x\in\mathcal{C}^{\prime}$. Then, by definition, $H_{1}$ is well-defined, and the same follows for each $H_{i}$ inductively. We conclude that $H$ is well-defined, as it is piecewise equivalent to the $H_{i}$’s.
To show (b), note that by Lemma D.2, $\mathcal{C}^{\prime}_{i}(x_{-i})$ is continuous in $x_{-i}$. Then, continuity of each $h_{i}$ in $(t,x)$ follows together with the continuity of the centroid function $\overline{c}(\cdot)$. Similarly as in (a), the recursive definitions of each $H_{i}$ and the piecewise definition of $H$ lets us conclude that $H$ is indeed continuous in $(t,x)$. In particular, we point out that continuity at the boundary of the pieces is ensured, as for all $i\in[m]\setminus\left\{1\right\}$ and all $x\in\mathcal{C}^{\prime}$, by definition, $\lim_{t\to{(i-1)/m}^{+}}H_{i}(t,x)=H_{i-1}(\varphi^{-1}_{i-1}(1),x)$.
@@
By a variant of Berge’s maximum theorem for strongly concave objective functions over convex constraints (see Theorem C.3 in Appendix C), $\Psi_{i}$ is continuous and single-valued. Define the joint best response map $\Psi:\mathcal{C}^{\prime}\to\mathcal{C}^{\prime}$ as the composition $\Psi\coloneqq\Psi_{1}\circ\dots\circ\Psi_{m}$, which is also continuous.
@@ E.2 Technical Lemma
We provide a lemma which is applied in the proof of Lemma F.8 in Appendix F. Informally, it says the following: If we perform a projected gradient ascent step along a (locally) smooth function $f$ and the observed gradients of $f$ are perturbed by some error $\delta$, then, as long as $\delta$ is sufficiently small, the value of $f$ will not decrease.
Lemma E.3.
@@
We restate the following sufficient increase inequality for projected gradient ascent that will be used in the proof of Theorem 4.6. The claimed inequality is standard; for reference, see Bubeck et al. (2015), Lemma 3.6.
Lemma E.4.
@@ F.1 Mixed-Strategy Potential Games
In Example 4.2 we claimed that when considering mixed strategies over a finite action potential game, the potential structure is lifted to the continuous action game over compact convex set of mixed strategies. Here we state the corresponding proposition.
Proposition F.1.
@@
For the case of unconstrained potential games, Proposition F.1 is a direct consequence of Monderer & Shapley (1996), Lemma 2.10. The same reasoning extends straightforwardly to the constrained case; we therefore omit the proof here.
@@ F.2 Technical Lemmas
Lemma F.3.
Let $\Phi:\mathcal{X}\to\mathbb{R}$ be a potential function for our game with utilities $\{u_{i}\}_{i\in[m]}$. Then $\Phi$ is $(M\sqrt{m})$-smooth in the joint strategy $x\in\mathcal{X}$, where $M$ is the smoothness parameter of the utilities, see Assumption 4.4.
@@
Note that differentiability of $u_{i}$ for all $i\in[m]$ (Assumption 4.4) implies differentiability of $\Phi$, and it is well known that for all $x\in\mathcal{X}$, $\nabla_{x_{i}}\Phi(x)=\nabla_{x_{i}}u_{i}(x)$. Using the smoothness of $u_{i}$ from Assumption 4.4, we get
@@ F.3 Smoothness Along The Trajectory
where we use $L$-Lipschitz continuity of $u_{i}$ and $c_{j}$, see Assumption 4.4, for bounding the gradient norms. Then, we have for each $j\in[b]$,
@@
Using $M$-smoothness and $L$-Lipschitz continuity of $c_{j}$ for all $j\in[b]$, as well as $(M\sqrt{m})$-smoothness of $\Phi$, as shown in Lemma F.3, we then bound the Hessian norm as follows:
@@
Lemma F.7.
Suppose $x^{(t)}\in\mathcal{C}$ is $\beta$-feasible and each $\gamma^{(t)}$ is bounded as in the statement of Lemma F.5. Then $\Phi^{\eta}$ is $M^{\eta,\beta/2}_{\Phi}$-smooth over $x^{(t)},x^{(t+1)}]$ where $M^{\eta,\beta}_{\Phi}$ is defined as in Lemma F.6, i.e. for all $x,x^{\prime}\in[x^{(t)},x^{(t+1)}]$,
@@
Taking the norm, and using the fact that $x^{\prime}+t(x-x^{\prime})$ is $\beta/2$-feasible for all $t\in[0,1]$ as shown in Lemma F.5, together with the Hessian norm bound from Lemma F.6, we get
@@ F.4 Lower Bounding The Stepsize
Motivated by the bound in Lemma F.5 and smoothness along the trajectory as shown in Lemma F.7, we choose the stepsize at iteration $t$ as
@@
In this subsection, we want to argue that the above defined stepsizes can be uniformly lower bounded by a constant multiple of $\eta$. This is necessary to guarantee sufficient progress in our convergence analysis. We start by showing that feasibility margins do not decrease further once a point sufficiently close to the boundary is reached. Based on Lemma E.3, we prove the following statement. Note that a similar statement was proven in Usmanova et al. (2024), Fact 1, with the difference that ours is for projected gradient steps, which requires a different approach.
@@
Lemma F.8.
Let Assumption 4.5 hold with $\rho\geq\eta$. If at any iteration $0\leq t\leq T-1$, we have $\beta^{(t)}\leq\overline{c}\eta$ with $\overline{c}=\frac{\ell}{6L(b+1)}$, then for iteration $t+1$, we get $\prod_{j\in\mathcal{J}}\beta_{j}^{(t+1)}\geq\prod_{j\in\mathcal{J}}\beta_{j}^{(t)}$ for any $\mathcal{J}$ such that $\mathcal{J}_{t}\subseteq\mathcal{J}$ with $\mathcal{J}_{t}\coloneqq\{j\in[b]\mid\beta_{j}^{(t)}\leq\eta\}$.
@@
We aim to apply Lemma E.3 with $x^{(t)}, x+=x$$^{+}=x^{(t+1)}$,
@@
Note that by our choice of $\gamma^{(t)}$, by Lemma F.7, we have $\lVert\nabla f(x^{+})-\nabla f(x)\rVert\leq\frac{1}{\gamma^{(t)}}\lVert x^{+}-x\rVert$. In order to apply Lemma E.3, we further need to show that for $F(x)\coloneqq\frac{1}{\gamma^{(t)}}\left(x^{+}-x\right)$, it holds that $\lVert F(x)\rVert\geq 3\lVert\delta\rVert$.
We begin by giving a lower bound on $\lVert F(x^{(t)})\rVert$. By Proposition B.1 of Agarwal et al. (2021) and Assumption 4.5, we have
@@
where $(\star)$ is due to $\beta_{j}^{(t)}>\eta$ for $j\not\in\mathcal{J}_{t}$. Combining the two bounds, and by definition of $\overline{c}$, we obtain from Lemma E.3 that
@@
As a consequence of Lemma F.8, we obtain the following lower bound on the distance of iterates to the boundary of $\mathcal{C}$. Having established Lemma F.8 above, the proof of Lemma F.9 is equivalent to the proof of an analogous result stated as Lemma 6 in Usmanova et al. (2024); we omit the proof here.
Lemma F.9.
Let Assumptions 4.3 and 4.5 hold with $\rho\geq\eta$. Then, for all $t\geq 1$, it holds that $\beta^{(t)}\geq c\eta$ where $c\coloneqq\frac{1}{2}\left(\frac{\ell}{24L(b+1)}\right)^{b}$.
@@
First, we argue that our independent updates are equivalent to centralized gradient ascent on the regularized potential function $\Phi^{\eta}$. This follows from the fact that, by Lemma F.2, the regularized utilities form a potential game, meaning we have $\nabla_{x_{i}}\Phi^{\eta}(x)=\nabla_{x_{i}}B^{\eta}_{i}(x)$ for any $x\in\mathcal{X}$ and $i\in[m]$. Together with separability of the projection operator onto the Cartesian product of the sets $\mathcal{X}_{i}$, see Leonardos et al. (2022), Lemma D.1, we obtain
@@
By smoothness along the trajectory and our choice of stepsizes, by Lemma E.4, for all $0\leq t\leq T-1$,
@@
Together with the fact that there exists a constant $C>0$ such that $\gamma^{(t)}\geq C\eta$, see Lemma F.10, and by telescoping the sum, we obtain
@@
We note that using Lemma F.9, $\Phi^{\eta}$ can be bounded by a constant. Thus for $\eta=\epsilon$ and $T=\mathcal{O}(\epsilon^{-3})$, there exists $0\leq t^{\star}\leq T-1$ such that $\lVert x^{(t^{\star}+1)}-x^{(t^{\star})}\rVert\leq\frac{\epsilon}{4}$. Next, we argue that then $x^{(t^{\star}+1)}$ satisfies the $\frac{\epsilon}{2}$-KKT conditions as introduced in Appendix E. Consider the pair of primal and dual variables given by
@@
    Primal feasibility: Lemma F.5 implies that all iterates remain within $\mathcal{C}$. In particular, we have $x^{\star}\in\mathcal{C}$.
@@
    Lagrangian stationarity: By a standard property of projected gradient ascent, see e.g. Agarwal et al. (2021), Proposition B.1, $\lVert x^{(t^{\star}+1)}-x^{(t^{\star})}\rVert\leq\frac{\epsilon}{4}$ implies that
@@
Finally, Lemma E.2 implies that the $\frac{\epsilon}{2}$-KKT strategy $x^{\star}$ constitutes an $\epsilon$-approximate constrained Nash equilibrium. ∎